### 12 Dec 2023 v2

# Parametrically enhancing sensor sensitivity at an exceptional point

Abstract
We propose a scheme to enhance the sensitivity of Non-Hermitian optomechanical mass-sensors. The benchmark system consists of two coupled optomechanical systems where the mechanical resonators are mechanically coupled. The optical cavities are driven either by a blue or red detuned laser to produce gain and loss, respectively. Moreover, the mechanical resonators are parametrically driven through the modulation of their spring constant. For a specific strength of the optical driving field and without parametric driving, the system features an Exceptional Point (EP). Any perturbation to the mechanical frequency (dissipation) induces a splitting (shifting) of the EP, which scales as the square root of the perturbation strength, resulting in a sensitivity-factor enhancement compared with conventional optomechanical sensors. The sensitivity enhancement induced by the shifting scenario is weak as compared to the one based on the splitting phenomenon. By switching on parametric driving, the sensitivity of both sensing schemes is greatly improved, yielding to a better performance of the sensor. We have also confirmed these results through an analysis of the output spectra and the transmissions of the optical cavities. In addition to enhancing EP sensitivity, our scheme also reveals nonlinear effects on sensing under splitting and shifting scenarios. This work sheds light on new mechanisms of enhancing the sensitivity of Non-Hermitian mass sensors, paving a way to improve sensors performance for better nanoparticles or pollutants detection, and for water treatment.
Optomechanics, exceptional point, sensor
pacs:
42.50.Wk, 42.50.Lc, 05.45.Xt, 05.45.Gg
December 12, 2023
I Introduction
Exceptional Points (EPs), Non-Hermitian degeneracy, are known in physical systems for their interesting counter-intuitive features and intriguing effects. Among the fascinating properties of EPs, we can mention nonreciprocity Thomas et al. (2016); Lau and Clerk (2018), topological transport phenomena Xu et al. (2023a, b), stopping light Goldzak et al. (2018), loss-induced suppression and revival of lasing, pump-induced lasing death and unidirectional invisibility (see Peng et al. (2014) and references therein). Recently, EPs have been used to enhance sensor sensitivity and versatile physical systems were proposed for this purpose including optical/photonic Hodaei et al. (2017); Chen et al. (2017), microwave Nada et al. (2018), plasmonics Park et al. (2020); Jiang et al. (2022), and mechanic/acoustic Rosa et al. (2021); Cai et al. (2022). Coupled optomechanical systems, which result from a coupling between an electromagnetic field and a mechanical object Djorwé et al. (2013a); Aspelmeyer et al. (2014), have also been used to engineer EPs owing to their ability to induce gain and losses depending on their sideband pumping mechanism. Exceptional points engineering in optomechanics have led to a number of interesting phenomena such as mass sensing improvement Djorwe et al. (2019a); Tchounda et al. (2023), topological energy transfer Chen et al. (2023); Xu et al. (2016), and nonlinearly induced collective phenomena Djorwe et al. (2018); Djorwé et al. (2020) among others.
Exceptional point sensing is an approach which provides better sensitivity compared to conventional sensing schemes where the perturbation induces a shift in the frequency. The resulting shift/signal is generally weak and needs a pre-amplification process in order to be detected and analyzed Zhu et al. (2021). For the Non-Hermitian sensors, however, the signal is self-amplified owing to the abrupt singularity and the topological feature at the EP, leading to a giant sensitivity enhancement without a pre-amplification requirement of the detected signal. This new sensing scheme has recently attracted attention as a means to enhance the responsiveness of sensors, and has started to be practically implemented. For instance, exceptional point sensing has been used to monitor internal physiological states via wireless interrogation techniques Chen and El-Ganainy (2019), and for an efficient wireless power transfer Assawaworrarit et al. (2017); Song et al. (2021). Beside these interesting sensing applications probed at EPs, the implementation of EP-sensors is still in its infancy stage and it seems to rise up scepticism regarding noise amplification, which induces a degradation of the sensor’s performance Anderson et al. (2023). This noise issue can be of a fundamental nature owing to the eigenbasis collapse or of a technical nature related to the amplification mechanisms used for the realization of EPs. This question has been a bottleneck of EP-sensors, however, appropriate techniques have started to be used to handle and suppress this noise issue Kononchuk et al. (2022); Suntharalingam et al. (2023); Tuxbury et al. (2022). Therefore, tremendous research activities are still going on to tackle and to get rid of this limitation Chen and Zhao (2020); Liu et al. (2021); Wu et al. (2023); Khanbekyan (2023), and this includes nonlinear phenomena Li et al. (2023); Bai et al. (2022).
The current work introduces a new mechanism to enhance the sensitivity of EP-sensors, which is based on a mechanical parametric driving effect. Our proposal consists of two coupled optomechanical systems, where the mechanical resonators are mechanically coupled. The parametric driving comes from a nonlinear modulation of the spring constants of the mechanical resonators Lemonde et al. (2016); Bothner et al. (2020). We investigated two sensing schemes. The first one is based on the splitting Djorwe et al. (2019a) of the EP and the second is related to the shifting Wang et al. (2019); Mao et al. (2023) of the EP. When parametric driving is switched on, we found that i) the sensitivity is greatly enhanced, ii) the EP is shifted towards low frequencies, and iii) the strength of the optical driving field corresponding to the EP increases. Moreover, it has been observed that the overall sensitivity enhancement in the shifting scenario is weak compared to the enhancement captured with the splitting scheme. Furthermore, the output spectra and transmissions of the optical cavities have been investigated under some system parameters. The effect of perturbations, the parametric driving impact and thermal noise on the EP were analyzed via the spectra. While the effect of the other parameters is confirmed as aforementioned, it has been found that thermal noise amplifies the output signal, leading to better sensitivity. This work shows how sensing based on the splitting of the EP outperforms sensors operating in a scenario where the EP is shifted. Furthermore, it also paves a way towards improvement of sensing through nonlinear effects in optomechanics Djorwé et al. (2013b); Leijssen et al. (2017); Djorwe et al. (2019b).
The rest of the work is organized as follows. In Section II we describe the model and derive the related dynamical equations. The sensitivity schemes based on splitting and shifting of the EP are presented in Section III. The transmissions and output spectra of the optical cavities are analyzed in Section IV, while Section V concludes our work.
II Model and dynamical equations
We propose a two-cavity system with mechanically coupled vibrating mechanical resonators, as shown in /cref[fig:Fig1]. One cavity is driven by a red (blue) driving field that creates loss (gain) in the system. In addition, the mechanical resonators are parametrically driven for sensitivity enhancement prospect as it will be pointed out later on. The Hamiltonian of this system ($\hbar=1$) is given by,
Figure 1: Sketch of our benchmark system. Cavity $1$ ($2$) is driven with red (blue) electromagnetic field to generate losses (gain). The sprint constants of the involved mechanical resonators are modulated to induce parametric driving Lemonde et al. (2016); Bothner et al. (2020).
$H=H_{\rm{0M}}+H_{\rm{PA}}+H_{\rm{int}}+H_{\rm{drive}},$ (1)
where,
$H_{\rm{0M}}$ $=$ $\sum_{j}\omega_{j}b^{\dagger}_{j}b_{j}+\omega_{c}^{j}a^{\dagger}_{j}a_{j}-g_{j}a^{\dagger}_{j}a_{j}(b^{\dagger}_{j}+b_{j}),$ (2)
$H_{\rm{PA}}$ $=$ $\sum_{j}\frac{\chi_{j}}{2}(b^{\dagger 2}_{j}e^{-i\phi_{d}}+b^{2}_{j}e^{-i\phi_{d}})$ (3)
$H_{\rm{int}}$ $=$ $-J_{m}(b_{1}b_{2}^{\dagger}+b_{1}^{\dagger}b_{2}),$ (4)
$H_{\rm{drive}}$ $=$ $i\sum_{j}(\rm{E_{j}}a^{\dagger}_{j}e^{-i\omega^{j}_{p}t}-\rm{E^{*}_{j}}a_{j}e^{i\omega^{j}_{p}t}).$ (5)
In Eq. 1, $H_{\rm{0M}}$, $H_{\rm{PA}}$, $H_{\rm{int}}$, and $H_{\rm{drive}}$ are the Hamiltonians describing the optomechanical cavities, the mechanical parametric amplification (MPA), the mechanical interactions and the driving fields, respectively. The annihilation (creation) operators related to the mechanical resonators and optical fields are $b_{j}$ ($b^{\dagger}_{j}$) and $a_{j}$ ($a_{j}^{\dagger}$), respectively. The other parameters are the mechanical ($\omega_{j}$) and cavity ($\omega_{c}^{j}$) frequencies, the optomechanical coupling ($g_{j}$), the phonon hopping coupling rate ($J_{m}$) and the driving amplitude ($\rm{E_{j}}$). The parametric driving amplitude $\chi_{j}$ comes from the modulation of the spring constants of the mechanical resonators at frequency $2\omega_{d}$ and phase $\phi_{d}$ Lemonde et al. (2016); Bothner et al. (2020). For the sake of simplicity and without loss of generality, we assume that the optical cavities are driven by the same amount of driving strength and are degenerated, so that from now on we consider $\rm{E_{1}}=\rm{E_{2}}=\rm{E}$, $\omega^{1}_{c}=\omega^{2}_{c}=\omega_{c}$, and $g_{1}=g_{2}=g$. Throughout the work, we also assume $\phi_{d}=0$ and $\chi_{1}=\chi_{2}=\chi$. In the rotating wave approximation (RWA) and in the frame rotating at $\omega_{p}^{j}+\omega_{d}$, where $\omega_{p}^{j}$ is the $j^{th}$ electromagnetic driving frequency, the above Hamiltonian yields:
$H_{\rm{0M}}$ $=\sum_{j}\tilde{\omega}_{j}b^{\dagger}_{j}b_{j}-\Delta_{j}a^{\dagger}_{j}a_{j}-ga^{\dagger}_{j}a_{j}(b^{\dagger}_{j}+b_{j})$ (6a)
$H_{\rm{PA}}$ $=\frac{\chi}{2}(b^{\dagger 2}_{j}+b^{2}_{j}),$ (6b)
$H_{\rm{int}}$ $=-J_{m}(b_{1}b_{2}^{\dagger}+b_{1}^{\dagger}b_{2}),$ (6c)
$H_{\rm{drive}}$ $=i\sum_{j}(\rm{E}a^{\dagger}_{j}-\rm{E}^{*}a_{j}),$ (6d)
where $\tilde{\omega}_{j}=\omega_{j}-\omega_{d}$ and $\Delta_{j}=\omega_{p}-\omega_{c}$. In order to diagonalize the quadratic term in the above Hamiltonian, we introduce the squeezing transformation $S(r_{j})=\rm{exp}[r_{j}(b^{2}_{j}-b^{\dagger 2}_{j})]$ where the squeezing parameter is defined as $r_{j}=\frac{1}{4}\log{\frac{\tilde{\omega}_{j}+\chi}{\omega_{j}-\chi}}$. By using the Bogoliubov transformation $b_{s_{j}}=b_{j}\cosh(r_{j})+b^{\dagger}_{j}\sinh(r_{j})$, and invoking the Heisenberg equation, one gets the following Quantum Langeving Equations (QLEs),
$\dot{a_{j}}$ $=(i[\Delta_{j}+\tilde{g}(b_{s_{j}}+b^{\dagger}_{s_{j}})]-\frac{\kappa}{2})a_{j}+\sqrt{\kappa}\alpha^{in}_{j}+\sqrt{\kappa}a^{in}_{j},$ (7a)
$\dot{b_{s_{j}}}$ $=-(i\Delta_{m}^{j}+\frac{\gamma_{m}}{2})b_{s_{j}}+i\tilde{J}_{m}b_{3-j}+i\tilde{g}a^{\dagger}_{j}a_{j}+\sqrt{\gamma_{m}}\beta^{in}_{j}.$ (7b)
In Section II, the driving amplitude $\rm{E}$ has been replaced by $\sqrt{\kappa}\alpha^{in}$ with $\alpha^{in}=\sqrt{\frac{P_{in}}{\hbar\omega_{p}}}$, where $P_{in}$ is the input power. The optical and mechanical dissipation rates are captured by $\kappa_{1}=\kappa_{2}=\kappa$ and $\gamma_{1}=\gamma_{2}=\gamma_{m}$, respectively. Here $a^{in}_{j}$ and $\beta^{in}_{j}$ are input noise operators of the j-th cavity and mechanical modes with zero mean values.
The nonlinear set of equations displayed in Section II can be linearized by splitting the operators in terms of their mean values ($c-$numbers) and some amount of quantum fluctuations as $a_{j}=\alpha_{j}+\delta\alpha_{j}$ and $b_{s_{j}}=\beta_{j}+\delta\beta_{j}$, where $\alpha_{j}=\langle a_{j}\rangle$ and $\beta_{j}=\langle b_{s_{j}}\rangle$. This procedure leads to classical mean equations,
$\dot{\alpha_{j}}$ $=(i\tilde{\Delta}_{j}-\frac{\kappa}{2})\alpha_{j}+\sqrt{\kappa}\alpha^{in},$ (8a)
$\dot{\beta_{j}}$ $=-(i\Delta_{m}^{j}+\frac{\gamma_{m}}{2})\beta_{j}+i\tilde{J}_{m}\beta_{3-j}+i\tilde{g}|\alpha_{j}|^{2},$ (8b)
and to the fluctuations equations,
$\delta\dot{\alpha_{j}}$ $=(i\tilde{\Delta}_{j}-\frac{\kappa}{2})\delta\alpha_{j}+i\tilde{G_{j}}(\delta\beta_{j}+\delta\beta^{\dagger}_{j})+\sqrt{\kappa}\delta\alpha^{in}_{j},$ (9a)
$\delta\dot{\beta_{j}}$ $=-(i\Delta_{m}^{j}+\frac{\gamma_{m}}{2})\delta\beta_{j}+i\rm{\tilde{J}_{m}\delta\beta_{3-j}}+i\tilde{G_{j}}(\delta\alpha^{\dagger}_{j}+\delta\alpha_{j})+\sqrt{\gamma_{m}}\delta\beta^{in}_{j},$ (9b)
where $\tilde{\Delta}_{j}=\Delta_{j}+2\tilde{g}\Re{\beta_{j}}$, $\tilde{G_{j}}=\tilde{g}\alpha_{j}$, and we have assumed $\alpha_{j}$ a real number.
In order to get an effective mechanical model of our proposal based on Section II, we need to trace out the optical equations. For that purpose, we introduce the following slowly varying operators with tildes, $\delta\alpha_{j}=\delta\tilde{\alpha_{j}}e^{i\tilde{\Delta}_{j}t}$, $\delta\alpha^{in}=\delta\tilde{\alpha^{in}}e^{i\tilde{\Delta}_{j}t}$, and $\delta\beta_{j}=\delta\tilde{\beta_{j}}e^{-i\Delta_{m}^{j}t}$. By using these slowly varying operators in Section II, and by neglecting the terms oscillating at higher frequencies, the set of dynamical equations for the fluctuations yields,
$\dot{\delta\tilde{\alpha_{1}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{1}}+i\tilde{G_{1}}\delta\tilde{\beta_{1}}+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{1},$ (10a)
$\dot{\delta\tilde{\alpha_{2}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{2}}+i\tilde{G_{2}}\delta\tilde{\beta_{2}}^{\dagger}+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{2},$ (10b)
$\dot{\delta\tilde{\beta_{1}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{1}}+i\tilde{J}_{m}\delta\tilde{\beta}_{2}+i\tilde{G_{1}}\delta\tilde{\alpha_{1}}+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{1},$ (10c)
$\dot{\delta\tilde{\beta_{2}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{2}}+i\tilde{J}_{m}\delta\tilde{\beta}_{1}+i\tilde{G_{2}}\delta\tilde{\alpha_{2}}^{\dagger}+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{2}.$ (10d)
The intracavity fields ($\delta\tilde{\alpha_{j}}$) in Section II can be integrated out (see Appendix A), and after some arrangements one obtains the following effective mechanical system,
$\dot{\delta\beta_{1}}=-(i\Delta_{m}^{1}+\frac{\rm{\Gamma_{eff}}^{1}}{2})\delta\beta_{1}+i\rm{\tilde{J}_{m}\delta\beta_{2}}+i\sqrt{\Gamma_{1}}\delta\alpha^{in}_{1}+\sqrt{\gamma_{m}}\delta\beta^{in}_{1},$ (11a)
$\dot{\delta\beta_{2}}=-(i\Delta_{m}^{2}+\frac{\rm{\Gamma_{eff}}^{2}}{2})\delta\beta_{2}+i\rm{\tilde{J}_{m}\delta\beta_{1}}+i\sqrt{\Gamma_{2}}\delta\alpha^{in\dagger}_{2}+\sqrt{\gamma_{m}}\delta\beta^{in}_{2},$ (11b)
where the optically induced damping is $\rm{\Gamma}_{j}=\frac{4|\tilde{G_{j}}|^{2}}{\kappa}$ and the effective damping are $\rm{\Gamma_{eff}}^{1,2}=\gamma_{m}\pm\rm{\Gamma}_{1,2}$. From now on, we will assume that $\rm{\Gamma}_{1}=\rm{\Gamma}_{2}=\rm{\Gamma}$. Section II can be put in its compact form as,
$\dot{\mathcal{O}}=\mathcal{M}\mathcal{O}+i\sqrt{\Gamma}\mathcal{N}+\sqrt{\gamma_{m}}\mathcal{O}^{in},$ (12)
where $\mathcal{O}\equiv(\delta\beta_{1},\delta\beta_{2})^{T}$, $\mathcal{N}\equiv(\delta\alpha^{in}_{1},\delta\alpha^{in\dagger}_{2})^{T}$, $\mathcal{O}^{in}\equiv(\delta\beta^{in}_{1},\delta\beta^{in}_{2})^{T}$, and the matrix
$\mathcal{M}=\begin{bmatrix}-(i\Delta_{m}^{1}+\frac{\rm{\Gamma_{eff}}^{1}}{2})&i\tilde{J}_{m}\\ i\tilde{J}_{m}&-(i\Delta_{m}^{2}+\frac{\rm{\Gamma_{eff}}^{2}}{2})\end{bmatrix}.$ (13)
III Sensitivity through a splitting and shifting of an exceptional point
In order to get the eigenvalues of our effective mechanical system, we rewrite Eq. 12 in the Schrodinger-like equation, $i\partial_{t}\Psi=\rm{H_{eff}}\Psi$ where $\Psi\equiv(\delta\beta_{1},\delta\beta_{2})^{T}$. By ignoring noise terms, this leads to an effective Hamiltonian,
$\rm{H_{eff}}=\begin{bmatrix}\Delta_{m}^{1}-i\frac{\rm{\Gamma_{eff}}^{1}}{2}&-\tilde{J}_{m}\\ -\tilde{J}_{m}&\Delta_{m}^{2}-i\frac{\rm{\Gamma_{eff}}^{2}}{2}\end{bmatrix}.$ (14)
Figure 2: Eigenmodes and the EP splitting (a,b) and shifting (c,d) after a perturbation strength of $\epsilon=5\times 10^{-3}\omega_{m}$. There is no parametric drive in both cases ($\chi=0$), and the $\rm{EP}$ happens around the driving strength $\alpha^{in}=420\omega_{m}^{-1/2}$ for both sensors when there is no perturbation. The dimensionless used parameters are, $\kappa=0.1\omega_{m}$, $\Delta_{1}=-\omega_{m}$, $\Delta_{2}=\omega_{m}$, $\gamma_{m}=10^{-3}\omega_{m}$, $g=2.5\times 10^{-4}\omega_{m}$, and $J_{m}=2.2\times 10^{-2}\omega_{m}$.
The eigenvalues $\lambda$ of $\rm{H_{eff}}$ are solutions of the following characteristic equation, $\det(\rm{H_{eff}}-\lambda\mathbb{I}_{2})$. Solving this equation leads to the eigenvalues,
$\lambda_{\pm}=\frac{1}{2}(\Delta_{m}^{1}+\Delta_{m}^{2})-\frac{i}{4}(\rm{\Gamma_{eff}}^{1}+\rm{\Gamma_{eff}}^{2})\pm\frac{\sigma}{4},$ (15)
where $\sigma=\sqrt{16\tilde{J}_{m}^{2}+(2\delta{\Delta_{m}}-i\delta\rm{\Gamma_{eff}})^{2}}$ with $\delta{\Delta_{m}}=\Delta_{m}^{1}-\Delta_{m}^{2}$ and $\delta\rm{\Gamma_{eff}}=\rm{\Gamma_{eff}}^{1}-\rm{\Gamma_{eff}}^{2}$. The eigenvalues of our system are given by Eq. 15, and are displayed in Fig. 2, where the eigenfrequencies ($\omega_{\pm}=\operatorname{Re}(\lambda_{\pm})$) and the dissipations ($\gamma_{\pm}=\operatorname{Im}(\lambda_{\pm})$) are shown in Fig. 2(a,c) and Fig. 2(b,d), respectively. In this figure, the parametric driving is off ($\chi=0$) and the solid (dashed) line depicts the case before (after) the perturbation of the system. In Fig. 2(a,b) the perturbation induces a splitting of the $\rm{EP}$ while in Fig. 2(c,d), it instead induces a shift of the $\rm{EP}$. In the former case, any mass deposition ($\delta m$) acts as a perturbation on the system, and induces a frequency shift ($\delta\omega$) through the relationship, $\delta m=\mathcal{R}^{-1}\delta\omega$ where $\mathcal{R}=\frac{\omega_{m}}{2m}$ is the mass responsibility and $\rm{m}$ is the mass of the resonator supporting the deposition. Therefore, this mass deposition leads to a frequency shift that lifts up the $\rm{EP}$-degeneracy, resulting in a sensitivity enhancement at the $\rm{EP}$ (see double arrow in Fig. 2(a,b)). In the latter case, the perturbation induces a shift in the dissipation instead, which can be seen as a dissipation mismatch ($\delta\gamma=|\gamma_{1}-\gamma_{2}|$) between the mechanical resonators Wang et al. (2019); Mao et al. (2023). This perturbation can also be thought of as induced by a mismatch between optical damping ($\delta\gamma=|\Gamma_{1}-\Gamma_{2}|$), and may originate from optical fluctuations within the cavities. This scenario also leads to a sensing scheme as shown by the double arrow in Fig. 2(c,d). A general observation from Fig. 2 is that after a splitting due to a mass deposition (Fig. 2(a,b)), the $\rm{EP}$ is destroyed and that limits the ability of the system to perform multiple sensing scenarios in a time sequence. Indeed, the enhanced performance is deteriorated after each sensing process since every next detection is conducted on the basis of being destroyed by the previous detection. In order to preserve $\rm{EP}$ in the parameter space after a perturbation, it is preferable to shift the $\rm{EP}$ instead of lifting it up. Such a sensor based on shifting the $\rm{EP}$ (see Fig. 2(c,d)) will therefore meet the condition of a non-demolition sensing mechanism. With a fine adjustment of gain and loss in the system, we expect a linewidth preservation of the eigenstates after optical perturbation in the sensing scheme based on shifting the $\rm{EP}$. This may improve the performance of such sensing scenarios in the perspective of experiments related to the precision of dispersion measurement Mao et al. (2023). However, we stress that the sensitivity of this novel sensing scheme is somehow weak compared to the splitting scheme (compare the double arrows in Fig. 2 (a) and (c)).
In order to enhance the sensitivity in the two aforementioned discussed sensing schemes, we turn on parametric driving ($\chi\neq 0$). Fig. 3 displays the enhancement of the splitting when parametric driving is accounted for in the system. Fig. 3a and Fig. 3c show the frequency splittings in the case of $\chi=0$ and $\chi=0.3\omega_{m}$, respectively. Their splitting ($\Delta\omega_{\pm}=\omega_{\pm}^{\epsilon}-\omega_{\pm}$) are displayed on Fig. 3b and Fig. 3d, respectively. It can be seen that the resulting splitting at the $\rm{EP}$ is greater in Fig. 3d than in Fig. 3b. This reveals the fact that parametric terms can be used to tune the performance of the sensor. Another observation is that, the parametric driving shifts the overall frequencies towards low frequencies (compare $y-axis$ of Fig. 3 (a) and (c)). The same investigation can also be done in sensing based on shifting, and similar observations can be pointed out (not shown).
Figure 3: Splitting of the eigenfrequencies for $\chi=0$ (a) and for $\chi=0.3\omega_{m}$ (c). The solid line is before any perturbation and the dashed line is after a mass deposition that has induced a shift of ($\delta\omega=2\times 10^{-3}\omega_{m}$ in both cases. (b) and (d) are the resulting splitting from (a) and (c), respectively. The used parameters are the same as in Fig. 2.
In order to get more insight into the performance of these sensing schemes, we plotted in Fig. 4 the splitting and the enhancement factor. Fig. 4(a,b) show these quantities for the sensing based on splitting while Fig. 4(c,d) display the same quantities for the sensing based on shifting effect. The splitting represented here is the quantity ($\Delta\omega_{\pm}=|\omega_{\pm}^{\epsilon}-\omega_{\pm}|$) at the $\rm{EP}$, while the enhancement factor is defined as being this splitting over the perturbation strength $\eta=|\frac{\Delta\omega_{\pm}}{\delta\omega}|$. In all these figures, the full line is without parametric driving ($\chi=0$), while the dashed-dotted and dashed lines are for $\chi=0.5\omega_{m}$ and $\chi=0.7\omega_{m}$, respectively. It can be seen that parametric driving improves both the sensitivity and the enhancement factor. Furthermore, one can observe that these quantities scale faster in the splitting sensing scheme as compared to shifting scenario sensing (compare Fig. 4a to Fig. 4c or Fig. 4b to Fig. 4d). One benefit of sensors operating under splitting mechanisms over those based on shifting is that the former ones perform very well for small perturbations, while the latter do not. Indeed, Fig. 4(a,b) show giant enhancement sensitivity for small perturbations, and that is owing to the square root topological feature at the $\rm{EP}$. However, such an enhancement for small perturbations is missing in Fig. 4(c,d), and is even zero (see inset). One reason behind this lack of sensitivity to small perturbations is that the shift is negligible for tiny disturbances
Figure 4: (a,c) Sensitivities versus the perturbation for different values of $\chi$. (b,d) Enhancement factors versus the perturbation for different values of $\chi$. (a,b) are from the splitting sensing scheme, while (c,d) result from the sensing approach based on shifting the $\rm{EP}$. The rest of the parameters are the same as in Fig. 2.
To understand the performance of these two sensing scenarios regarding the parametric term, one evaluates the quantity $\Delta\lambda_{\pm}=\lambda_{\pm}^{\epsilon}-\lambda_{\pm}$. This term is the difference between the perturbed eigenvalue $\lambda_{\pm}^{\epsilon}$ and the non-perturbed one, where $\epsilon$ is the strength of the perturbation. As stated above, our proposal can experience sensing through both splitting and shifting of the $\rm{EP}$. As we are interested in quantifying the sensitivity at the $\rm{EP}$, we can derive this eigenvalue’s difference at the $\rm{EP}$ which yields (see Appendix B),
$\Delta\lambda_{\pm}^{\rm{EP}}=\frac{\Delta\omega^{\epsilon}}{2}-\frac{i|\delta\gamma|}{4}\pm\frac{\sigma_{\rm{EP}}^{\epsilon}}{2},$ (16)
where
$\Delta\omega^{\epsilon}=\frac{1}{2}\left[\frac{\omega_{m}\delta\omega}{\Delta_{m}^{2}}-\frac{1}{2}\left(\frac{\chi_{1}^{2}}{\omega_{m}}+\frac{\chi_{2}^{2}}{\omega_{m}}\right)\right],$ (17)
and
$\sigma_{\rm{EP}}^{\epsilon}=\sqrt{(8J_{m}r)^{2}+4\nu^{2}-4i\nu|\delta\gamma|+4\nu\mu-2i\mu|\delta\gamma|},$ (18)
with $\nu=\frac{1}{2}\left(\frac{\chi_{2}^{2}}{\omega_{m}}-\frac{\chi_{1}^{2}}{\omega_{m}}\right)-\frac{\omega_{m}\delta\omega}{\Delta_{m}^{2}}$ and $\mu=-i\delta\Gamma_{\rm{eff}}$. In these equations (Eq. 17 and Eq. 18), all scenarios of perturbations allowed in our system have been considered. By considering that there is no parametric driving ($\chi_{j}=0$), and that there is no perturbation of the dissipation ($\delta\gamma=0$), the system reduce to the one investigated in Djorwe et al. (2019a), and the expressions in Eq. 17 and Eq. 18 reduce to $\Delta\omega^{\epsilon}=\frac{\delta\omega_{2}}{2}$ and $\sigma_{\rm{EP}}^{\epsilon}=(1+i)\sqrt{\frac{\delta\Gamma_{\rm{eff}}\delta\omega_{2}}{8}}$ as expected. This sensing scenario is based on a splitting at the $\rm{EP}$ as it can be seen in Fig. 2(a,b). In that case, the amplification factor is given by:
$\eta=\left|\frac{\Re{(\Delta\lambda_{\pm}^{\rm{EP}})}}{\delta\omega}\right|=\sqrt{\frac{\delta\Gamma_{\rm{eff}}}{8\delta\omega}}=\sqrt{\frac{m\delta\Gamma_{\rm{eff}}}{4\omega_{m}\delta m}}.$ (19)
When the perturbation is mainly from the dissipation ($\chi_{j}=0$, $\delta\omega=0$, $\delta\gamma\neq 0$), which can be caused either by the mismatch damping of the two mechanical resonators or the fluctuation of the gain and loss induced by a driving fluctuation, both sensitivity and amplification factor are reduced to,
$\Delta\lambda_{\pm}^{\rm{EP}}=-\frac{i}{4}|\delta\gamma|\pm\frac{i}{4}\sqrt{2|\delta\gamma|\delta\Gamma_{\rm{eff}}},$ (20)
and
$\eta=\left|\frac{\Re{(\Delta\lambda_{\pm}^{\rm{EP}})}}{\delta\omega}\right|=\sqrt{\frac{\delta\Gamma_{\rm{eff}}}{8|\delta\gamma|}},$ (21)
which also scales as a square root of the perturbation. It should be highlighted that $\delta\gamma$ can be negative in order to make sense of the enhancement factor definition. Moreover, Eq. 20 shows that there is no effect of the perturbation on the reference frequency while there is a negative shift of $\frac{1}{4}|\delta\gamma|$ in the reference dissipation as seen in Fig. 2(c,d). When parametric mechanical driving is involved ($\chi_{j}\neq 0$), we expect to point out some improvement as depicted in Fig. 4. By considering the sensing based on splitting scheme for instance ($\chi_{j}\neq 0$, $\delta\omega\neq 0$, $\delta\gamma=0$), one deduces from Eq. 16 the expression for the frequency difference at the $\rm{EP}$ as,
$\Delta\lambda_{\pm}^{\rm{EP}}=\frac{1}{2}\left(\frac{\omega_{m}}{\Delta_{m}^{2}}\delta\omega-\frac{\chi^{2}}{\omega_{m}}\right)\pm\frac{1}{2}\sqrt{(8J_{m}r)^{2}+i\frac{\omega_{m}}{\Delta_{m}^{2}}\delta\omega\delta\Gamma_{\rm{eff}}}.$ (22)
In this expression, it results that the parametric driving shifts the reference frequency through $\frac{\chi^{2}}{\omega_{m}}$ as it appears in Fig. 3c (compare $\rm{y-axis}$ of (a) and (c)), and also enhances the sensitivity. Indeed, the term ($(8J_{m}r)^{2}$) which comes from the expansion of $\rm{\cosh{2r}}$ not only shifts the $\rm{EP}$ from the driving strength $\alpha^{in}\sim 420\omega_{m}^{1/2}$ (see Fig. 3a) to $\alpha^{in}\sim 500\omega_{m}^{1/2}$ (see Fig. 3c), but also contributes to the enhancement of the sensitivity as depicted in Fig. 4(a,c). Furthermore, it should also be underlined that $\chi$ induces an amount of sensitivity improvement through the the term $\frac{\omega_{m}}{\Delta_{m}^{2}}$ under the square-root in Eq. 22. Similar explanations regarding the effects of parametric driving on the sensitivity can be done for the sensing scenario based on the shifting mechanism. In fact, the expression of the frequency difference at the $\rm{EP}$ in this case is derived as,
$\Delta\lambda_{\pm}^{\rm{EP}}=-\frac{\chi^{2}}{2\omega_{m}}-\frac{i}{4}|\delta\gamma|\pm\frac{1}{4}\sqrt{(8J_{m}r)^{2}-2|\delta\gamma|\delta\Gamma_{\rm{eff}}}.$ (23)
From this expression, we can see that $\chi$ induces a shift of the reference frequency through $\frac{\chi^{2}}{2\omega_{m}}$ as seen while comparing the $\rm{y-axis}$ of Fig. 3a to Fig. 3c. As in Fig. 2d, the shift in the dissipation is still captured by $\frac{|\delta\gamma|}{4}$. More importantly, in Eq. 23, the sensitivity is mainly enhanced by the parametric driving through $(8J_{m}r)^{2}$. Therefore, these analytical expressions predict the improvement of both sensing schemes exhibited by our proposal, and are in good agreement with our findings.
IV Transmissions and output spectra
The current section analyzes the transmission and output spectra of the cavity fields. The effects of thermal noise, and the different involved perturbations will be investigated in order to figure out their impact on the EP feature. For that purpose, we start from Section II (or its analogous without tilde), and by moving to the frequency domain we derive the following scattering coefficients (see details in Appendix C),
$S_{11}(\omega)$ $=1-\dfrac{2\tilde{G}_{1}\xi_{2}(\omega)\sqrt{\Gamma_{1}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{12}(\omega)=-\dfrac{2i\tilde{G}_{1}\tilde{J}_{m}\sqrt{\Gamma_{2}}}{\sqrt{\kappa}\xi_{eff}(\omega)},$ (24a)
$S_{13}(\omega)$ $=\dfrac{2i\tilde{G}_{1}\xi_{2}(\omega)\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{14}(\omega)=\dfrac{2\tilde{G}_{1}\tilde{J}_{m}\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)},$ (24b)
$S_{21}(\omega)$ $=\dfrac{2i\tilde{G}_{2}\tilde{J}_{m}\sqrt{\Gamma_{1}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{22}(\omega)=1+\dfrac{2\tilde{G}_{2}\xi_{1}(\omega)\sqrt{\Gamma_{2}}}{\sqrt{\kappa}\xi_{eff}(\omega)},$ (24c)
$S_{23}(\omega)$ $=\dfrac{2\tilde{G}_{2}\tilde{J}_{m}\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{24}(\omega)=-\dfrac{2i\tilde{G}_{2}\xi_{1}(\omega)\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)},$ (24d)
where $\xi_{eff}(\omega)=\xi_{1}(\omega)\xi_{2}(\omega)+\tilde{J}^{2}_{m}$ with $\xi_{1}(\omega)=\Gamma^{1}_{eff}/2+i(\Delta^{1}_{m}-\omega)$ and $\xi_{2}(\omega)=\Gamma^{2}_{eff}/2+i(\Delta^{2}_{m}-\omega)$. The coefficients $S_{11}$ and $S_{22}$ stand for the transmission of the first and second cavity, respectively.
Figure 5: (a,b) Transmission $S_{11}$ and $S_{22}$ of the first and second cavities, respectively. (c,d) Output spectra $S_{1}^{out}$ and $S_{2}^{out}$ of the first and second cavities, respectively. The full line depicts the case at the EP for $\alpha^{in}=420\omega_{m}^{1/2}$ (see Fig. 2 (a)). The dashed and dash-dotted lines correspond to cases away from the EP, and the other parameters are the same as in Fig. 2
Owing to our assumption that the cavities are degenerated, these two coefficients become similar when the driving is strong enough (see full lines in Fig. 5 (a) and (b)). For weak driving strength ($\alpha^{in}\lesssim 90\omega_{m}^{1/2}$), however, $S_{11}$ shows two dips while $S_{11}$ exhibits two peaks because they are far from the EP (not shown). This weak driving regime is interesting for nonreciprocal phenomena, and will be developed in our future investigations. The full lines in Fig. 5 show the shape of transmissions (Fig. 5 (a) and (b)) and output spectra (Fig. 5 (c) and (d)) at the EP. The output spectra can be expressed in terms of the given coefficients in Section IV and yield (see Appendix C),
$S^{(1)}_{out}(\omega)$ $=\left|S_{11}(\omega)\right|{}^{2}+\left|S_{12}(\omega)\right|{}^{2}+\left(2N_{s_{1}}+1\right)\left|S_{13}(\omega)\right|{}^{2}$
$+\left(2N_{s_{2}}+1\right)\left|S_{14}(\omega)\right|{}^{2},$ (25a)
$S^{(2)}_{out}(\omega)$ $=\left|S_{21}(\omega)\right|{}^{2}+\left|S_{22}(\omega)\right|{}^{2}+\left(2N_{s_{1}}+1\right)\left|S_{23}(\omega)\right|{}^{2}$
$+\left(2N_{s_{2}}+1\right)\left|S_{24}(\omega)\right|{}^{2}.$ (25b)
where $N_{s_{1,2}}=(n_{th_{1,2}}+1)\sinh^{2}(r_{1,2})+n_{th_{1,2}}\cosh^{2}(r_{1,2})$, and $n_{th}$ the thermal population. As it can be seen, the shapes of these spectra are mainly depending on those of the transmission ones. The dashed and dash-dotted lines in Fig. 5 depict the situations away from the EP, and we can see two peaks which are reminiscent of the eigenfrequencies shown in Fig. 2(a) for instance. In what follows, we will only consider the spectrum $S_{1}^{out}$ and will analyze the effects of some parameters on it as displayed in Fig. 6. The effect of thermal noise $n_{th}$ on the output spectrum is shown in Fig. 6(a), where we see that the EP is kept even though the output signal is amplified. This means that any output signal from our EP-sensor will be amplified, and this is due to the gain in our system. Such an amplification phenomenon is well-known in optomechanical systems, and seems to be beneficial for Non-Hermitian sensing. Moreover, this thermal noise effect can be monitored by tuning the squeezing parameter $r$ through $N_{s}$. Fig. 6(b) depicts the effect of parametric driving on the output spectrum, and it can be observed that $\chi$ shifts the spectrum’s shape towards low-frequency values as expected (see Fig. 3(c) for instance). Moreover, the driving strength at which EP happens has been increased. Owing to these changes in the input driving, the EPs corresponding to both the dashed and the dash-dotted lines are missing in Fig. 6(b). To capture these EPs and to quantify how they have been shifted, one needs to adjust the driving strength to $\alpha^{in}\sim 130\omega_{m}^{1/2}$ for the dashed line and $\alpha^{in}\sim 170\omega_{m}^{1/2}$ for the dash-dotted line. Fig. 6(c) and Fig. 6(d) display the effects of the perturbations on the frequency and the dissipation (or quality factor), respectively. The former corresponds to a sensing scheme based on the splitting of the EP, while the latter relies on sensing upon shifting of the EP. As aforementioned in Section III, it can be seen that the EP is destroyed in Fig. 6(c) while it is preserved and shifted (down) in Fig. 6(d). Moreover, the preserved peak of the EP in Fig. 6(d) has been de-amplified and we stress that this may impair the performance of the sensing process.
Figure 6: Dynamics of the EP in the output spectrum versus some system parameters. The effect of the thermal noise in (a), and the parametric driving in (b), the frequency shift perturbation in (c), and dissipation mismatch perturbation in (d). The full lines captures the EP at $\alpha^{in}=420\omega_{m}^{1/2}$ as in Fig. 2(a), while the dashed and dash-dotted lines correspond to the perturbed cases depending on the studied parameter. The other used parameters are the same as in Fig. 2.
V Conclusion
We investigated the effects of mechanical parametric driving on sensing performance at an exceptional point. This parametric behaviour comes from the modulation of the mechanical spring constant of the involved resonators. Two sensing schemes have been put forward. The sensing upon splitting of the EP as the frequency undergoes a perturbation, and the sensing through shifting of the EP that is induced by a damping mismatch for instance. It has been pointed out that mechanical parametric driving enhances the sensitivity of both schemes. Moreover, this parametric driving shifts the frequency at which EP happens towards low values. Furthermore, we figured out the fact that the sensitivity enhancement is more efficient under the splitting scenario than what is realized while shifting the EP. These results have also been confirmed through output spectra and transmission analysis of the optical cavities. These parametric effects on sensing reveal how nonlinearities can be used as requirement tools for better performances of sensors. This work provides a new way of improving EP-sensors and sheds light on the beneficial effects of nonlinear phenomena on Non-Hermitian sensing.
Acknowledgments
This work has been carried out under the Iso-Lomso Fellowship at Stellenbosch Institute for Advanced Study (STIAS), Wallenberg Research Centre at Stellenbosch University, Stellenbosch 7600, South Africa. P. Djorwe acknowledges the receipt of a grant from the APS-EPS-FECS-ICTP Travel Award Fellowship Programme (ATAP), Trieste, Italy. M. Asjad and D. Dutykh have been supported by the Khalifa University of Science and Technology under Award No. FSU- 2023-014.
Appendix A Effective mechanical Hamiltonian
Some details about derivation of the effective mechanical Hamiltonian are presented in this Appendix. In the rotating wave approximation (RWA) and in the frame rotating at $\omega_{p}^{j}+\omega_{d}$, where $\omega_{p}^{j}$ is the $j^{th}$ driving frequency, our system’s Hamiltonian yields:
$H_{\rm{0M}}$ $=\sum_{j}\tilde{\omega}_{j}b^{\dagger}_{j}b_{j}-\Delta_{j}a^{\dagger}_{j}a_{j}-ga^{\dagger}_{j}a_{j}(b^{\dagger}_{j}+b_{j})+\frac{\chi}{2}(b^{\dagger 2}_{j}+b^{2}_{j})$ (26a)
$H_{\rm{int}}$ $=-J_{m}(b_{1}b_{2}^{\dagger}+b_{1}^{\dagger}b_{2}),$ (26b)
$H_{\rm{drive}}$ $=iE\sum_{j}(a^{\dagger}_{j}-a_{j}),$ (26c)
where $\tilde{\omega}_{j}=\omega_{j}-\omega_{d}$ and $\Delta_{j}=\omega_{p}-\omega_{c}$. This Hamiltonian can be diagonalized by introducing the squeezing transformation $S(r)=\rm{exp}[r(b^{2}-b^{\dagger 2})]$, where $\rm{r}$ will be figured out latter on. This operator transforms through the Bogoliubov transformation $b$ to $b_{s}$ as,
$b_{s}=b\cosh(r)+b^{\dagger}\sinh(r).$ (27)
From Eq. 27, it is straightforward to get,
$b_{j}$ $=b_{s_{j}}\cosh(r_{j})-b^{\dagger}{s_{j}}\sinh(r_{j})$ (28a)
$b^{\dagger}_{j}$ $=b{\dagger}_{s_{j}}\cosh(r_{j})-b_{s_{j}}\sinh(r_{j}),$ (28b)
which can be used in Appendix A, and that leads to the following transformed Hamiltonian,
$H_{\rm{0M}}$ $=\sum_{j}\Delta_{m}^{j}b^{\dagger}_{s_{j}}b_{s_{j}}-\Delta_{j}a^{\dagger}_{j}a_{j}-\tilde{g}a^{\dagger}_{j}a_{j}(b^{\dagger}_{s_{j}}+b_{s_{j}})$ (29a)
$H_{\rm{int}}$ $=-\tilde{J}_{m}(b_{s_{1}}b_{s_{2}}^{\dagger}+b_{s_{1}}^{\dagger}b_{s_{2}}),$ (29b)
$H_{\rm{drive}}$ $=iE\sum_{j}(a^{\dagger}_{j}-a_{j}),$ (29c)
$H_{\rm{diss}}^{S}$ $=-i\frac{\kappa}{2}\sum_{j}a^{\dagger}_{j}a_{j}-i\frac{\gamma_{m}}{2}\sum_{j}b^{\dagger}_{s_{j}}b_{s_{j}},$ (29d)
where the Hamiltonian $H_{\rm{diss}}$ captures the optical ($\kappa$) and the mechanical ($\gamma_{1}=\gamma_{2}=\gamma_{m}$) dissipation in the system. Furthermore, the driving amplitude has been expressed as $E=\sqrt{\kappa}a^{in}$, where $\alpha^{in}$ is related to the input power $P_{in}$ through $\alpha^{in}=\sqrt{\frac{P_{in}}{\hbar\omega_{p}}}$. The following parameters have been also defined,
$\tilde{g}$ $=ge^{-r},$ (30a)
$\Delta_{m}^{j}$ $=\sqrt{\tilde{\omega}_{j}^{2}-\chi^{2}},$ (30b)
$\tilde{J}_{m}$ $=J_{m}\cosh(2r).$ (30c)
The Quantum Langeving Equations (QLEs) of the Hamiltonian given in Appendix A can be derived using the Heisenberg equation $\dot{\mathcal{O}}=i[H,\mathcal{O}]=\mathcal{N}$, where $H$ is a Hamiltonian, $\mathcal{O}$ is an operator ($\mathcal{O}\equiv b_{s_{j}},a_{j}$), and $\mathcal{N}$ the quantum noise related to the operator. Following this procedure, the QLEs of our system yield,
$\dot{a}_{j}$ $=$ $(i[\Delta_{j}+\tilde{g}(b_{s_{j}}+b^{\dagger}_{s_{j}})]-\frac{\kappa}{2})a_{j}+\sqrt{\kappa}\alpha^{in}_{j}+\sqrt{\kappa}a^{in}_{j},$ (31a)
$\dot{b}_{s_{j}}$ $=$ $-(i\Delta_{m}^{j}+\frac{\gamma_{m}}{2})b_{s_{j}}+i\tilde{J}_{m}b_{3-j}+i\tilde{g}a^{\dagger}_{j}a_{j}+\sqrt{\gamma_{m}}\beta^{in}_{j}.$ (31b)
The nonlinear set of equations displayed in Appendix A can be linearized by splitting the operators in term of their mean values ($c-$numbers) and some amount of quantum fluctuations as $a_{j}=\alpha_{j}+\delta\alpha_{j}$ and $b_{s_{j}}=\beta_{j}+\delta\beta_{j}$, where $\alpha_{j}=\langle a_{j}\rangle$ and $\beta_{j}=\langle b_{s_{j}}\rangle$. This procedure leads to the classical mean equations,
$\dot{\alpha_{j}}$ $=(i\tilde{\Delta}_{j}-\frac{\kappa}{2})\alpha_{j}+\sqrt{\kappa}\alpha^{in}_{j},$ (32a)
$\dot{\beta_{j}}$ $=-(i\Delta_{m}^{j}+\frac{\gamma_{m}}{2})\beta_{j}+i\tilde{J}_{m}\beta_{3-j}+i\tilde{g}|\alpha_{j}|^{2},$ (32b)
and to the fluctuations equations,
$\delta\dot{\alpha}_{j}$ $=(i\tilde{\Delta}_{j}-\frac{\kappa}{2})\delta\alpha_{j}+i\tilde{G_{j}}(\delta\beta_{j}+\delta\beta^{\dagger}_{j})+\sqrt{\kappa}\delta\alpha^{in}_{j},$ (33a)
$\rm{\delta\dot{\beta}_{j}}$ $=\rm{-(i\Delta_{m}^{j}+\frac{\gamma_{m}}{2})\delta\beta_{j}}\rm{+i\tilde{J}_{m}\delta\beta_{3-j}}\rm{+i\tilde{G_{j}}(\delta\alpha^{\dagger}_{j}+\delta\alpha_{j})}$
$+\rm{\sqrt{\gamma_{m}}\delta{\beta}^{in}_{j}},$ (33b)
with $\tilde{\Delta}_{j}=\Delta_{j}+2\tilde{g}\Re{\beta_{j}}$, and $\tilde{G_{j}}=\tilde{g}\alpha_{j}$ where the intracavity field amplitude $\alpha_{j}$ is assumed to be real. In order to get an effective mechanical model describing Appendix A, we need to trace out the optical equations. For that purpose, we introduce the following slowly varying operators with tildes, $\delta\alpha_{j}=\delta\tilde{\alpha_{j}}e^{i\tilde{\Delta}_{j}t}$, $\delta\alpha^{in}=\delta\tilde{\alpha^{in}}e^{i\tilde{\Delta}_{j}t}$, and $\delta\beta_{j}=\delta\tilde{\beta_{j}}e^{-i\Delta_{m}^{j}t}$. By using these slowly varying operators in Appendix A, one get
$\dot{\delta\tilde{\alpha_{j}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+i\tilde{G_{j}}(\delta\tilde{\beta_{j}}e^{-i(\Delta_{m}^{j}+\tilde{\Delta})t}+\delta\tilde{\beta_{j}}^{\dagger}e^{i(\Delta_{m}^{j}-\tilde{\Delta}_{j})t})$ (34a)
$+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{j},$ (34b)
$\dot{\delta\tilde{\beta_{j}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+i\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+i\tilde{G_{j}}(\delta\tilde{\alpha_{j}}^{\dagger}e^{i(\Delta_{m}^{j}-\tilde{\Delta}_{j})t}$ (34c)
$+\delta\tilde{\alpha_{j}}e^{i(\Delta_{m}^{j}+\tilde{\Delta}_{j})t})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{j}.$ (34d)
As gain and losses are involved in our system, we assume that the first cavity is driven at the red sideband ($\tilde{\Delta}_{1}=-\Delta_{m}^{j}$) while the second one is driven at the blue sideband ($\tilde{\Delta}_{2}=\Delta_{m}^{j}$). Therefore, on the red sideband we get ($j=1$),
$\delta\tilde{\dot{\alpha_{j}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+i\tilde{G_{j}}(\delta\tilde{\beta_{j}}+\delta\tilde{\beta_{j}}^{\dagger}e^{2i\Delta_{m}^{j}t})+\sqrt{\kappa}\delta\tilde{\alpha^{in}_{j}},$ (35a)
$\delta\tilde{\dot{\beta_{j}}}$ $=\rm{-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+i\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+i\tilde{G_{j}}(\delta\tilde{\alpha_{j}}^{\dagger}e^{2i\tilde{\Delta}_{j}t}+\delta\tilde{\alpha_{j}})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{j}},$ (35b)
while on the blue sideband we have ($j=2$),
$\delta\tilde{\dot{\alpha_{j}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+i\tilde{G_{j}}(\delta\tilde{\beta_{j}}e^{-2i\Delta_{m}^{j}t}+\delta\tilde{\beta_{j}}^{\dagger})+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{j},$ (36a)
$\delta\tilde{\dot{\beta_{j}}}$ $=\rm{-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+i\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+i(\tilde{G_{j}}\delta\tilde{\alpha_{j}}^{\dagger}+\tilde{G_{j}}^{\ast}\delta\tilde{\alpha_{j}}e^{2i\tilde{\Delta}_{j}t})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{j}}.$ (36b)
In the RWA, the terms oscillating with high frequency can be ignored and the set of dynamical equations describing the system reduces to,
$\dot{\delta\tilde{\alpha_{1}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{1}}+i\tilde{G_{1}}\delta\tilde{\beta_{1}}+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{1},$ (37a)
$\dot{\delta\tilde{\alpha_{2}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{2}}+i\tilde{G_{2}}\delta\tilde{\beta_{2}}^{\dagger}+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{2},$ (37b)
$\dot{\delta\tilde{\beta_{1}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{1}}+i\tilde{J}_{m}\delta\tilde{\beta}_{2}+i\tilde{G_{1}}^{\ast}\delta\tilde{\alpha_{1}}+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{1},$ (37c)
$\dot{\delta\tilde{\beta_{2}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{2}}+i\tilde{J}_{m}\delta\tilde{\beta}_{1}+i\tilde{G_{2}}\delta\tilde{\alpha_{2}}^{\dagger}+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{2}.$ (37d)
In order to get the effective mechanical system, we need to trace out the intracavity field ($\delta\tilde{\alpha_{j}}$) in Appendix A. For this purpose, let us introduce $\delta\tilde{\alpha_{j}}=\delta\alpha_{m}^{j}e^{-\frac{\kappa}{2}t}$, which after using them in Appendix A leads to,
$\dot{\delta\alpha_{m}^{1}}$ $=(i\tilde{G_{1}}\delta\tilde{\beta_{1}}+\sqrt{\kappa}\delta\tilde{\alpha^{in}_{1}})e^{\frac{\kappa}{2}t},$ (38a)
$\dot{\delta\alpha_{m}^{2}}$ $=(i\tilde{G_{2}}\delta\tilde{\beta_{2}}^{\dagger}+\sqrt{\kappa}\delta\tilde{\alpha^{in}_{2}})e^{\frac{\kappa}{2}t}.$ (38b)
These equations can be integrated,
[…]
where in the weak coupling regime ($\kappa\gg\tilde{G_{j}}$), can be adiabatically integrated and yields,
$\delta\tilde{\alpha}_{1}$ $=\frac{2}{\kappa}(i\tilde{G_{1}}\delta\tilde{\beta_{1}}+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{1}),$ (40a)
$\delta\tilde{\alpha}_{2}$ $=\frac{2}{\kappa}(i\tilde{G_{2}}\delta\tilde{\beta_{2}}^{\dagger}+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{2}).$ (40b)
By using Appendix A back in Appendix A, we obtain the following effective mechanical system,
$\dot{\delta\beta_{1}}$ $=-(i\Delta_{m}^{1}+\frac{\rm{\Gamma_{eff}}^{1}}{2})\delta\beta_{1}+i\tilde{J}_{m}\delta\beta_{2}+i\sqrt{\Gamma_{1}^{\ast}}\delta\alpha^{in}+\sqrt{\gamma_{m}}\delta\beta^{in}_{1},$ (41a)
$\dot{\delta\beta_{2}}$ $=-(i\Delta_{m}^{2}+\frac{\rm{\Gamma_{eff}}^{2}}{2})\delta\beta_{2}+i\tilde{J}_{m}\delta\beta_{1}+i\sqrt{\Gamma_{2}}\delta\alpha^{in\dagger}+\sqrt{\gamma_{m}}\delta\beta^{in}_{2},$ (41b)
where optical induced damping is $\rm{\Gamma}_{j}=\frac{4|\tilde{G_{j}}|^{2}}{\kappa}$ and the effective damping are $\rm{\Gamma_{eff}}^{1,2}=\gamma_{m}\pm\rm{\Gamma}_{1,2}$. From now on, we will assume that $\rm{\Gamma}_{1}=\rm{\Gamma}_{2}=\rm{\Gamma}$. Appendix A can be put in its compact form as,
$\dot{\mathcal{O}}=\mathcal{M}\mathcal{O}+i\sqrt{\mathcal{K}}\mathcal{N}+\sqrt{\gamma_{m}}\mathcal{O}^{in}$ (42)
where $\mathcal{O}\equiv(\delta\beta_{1},\delta\beta_{2})^{T}$, $\mathcal{K}\equiv(\Gamma_{1}^{\ast},\Gamma_{1})^{T}$, $\mathcal{N}\equiv(\delta\alpha^{in},\delta\alpha^{in\dagger})^{T}$ ,$\mathcal{O}^{in}\equiv(\delta\beta^{in}_{1},\delta\beta^{in}_{2})^{T}$, and the matrix
$\mathcal{M}=\begin{bmatrix}-(i\Delta_{m}^{1}+\frac{\rm{\Gamma_{eff}}^{1}}{2})&i\tilde{J}_{m}\\ i\tilde{J}_{m}&-(i\Delta_{m}^{2}+\frac{\rm{\Gamma_{eff}}^{2}}{2})\end{bmatrix}.$ (43)
Appendix B Splitting vs shifting sensing schemes
This Appendix is meant to provide a comparison between the two sensing schemes investigated in the main text. We have shown the possibility of sensing based either the splitting or the shifting of the $\rm{EP}$, but a straight comparison between the two is missing. A general eigenvalue difference evaluated at the $\rm{EP}$ can be given by,
$\Delta\lambda_{\pm}^{\rm{EP}}=\frac{\Delta\omega^{\epsilon}}{2}-\frac{i|\delta\gamma|}{4}\pm\frac{\sigma_{\rm{EP}}^{\epsilon}}{2},$ (44)
where
$\Delta\omega^{\epsilon}=\frac{1}{2}\left[\frac{\tilde{\omega}_{2}\delta\omega_{2}}{\Delta_{m}^{2}}-\frac{\chi_{2}\delta\chi_{2}}{\Delta_{m}^{2}}-\frac{\chi_{1}\delta\chi_{1}}{\Delta_{m}^{1}}-\frac{1}{2}\left(\frac{\chi_{1}^{2}}{\tilde{\omega}_{1}}+\frac{\chi_{2}^{2}}{\tilde{\omega}_{2}}\right)\right],$ (45)
and
$\sigma_{\rm{EP}}^{\epsilon}=\sqrt{(8J_{m}r)^{2}+4\nu^{2}-4i\nu|\delta\gamma|+4\nu\mu-2i\mu|\delta\gamma|},$ (46)
with $\nu=\frac{1}{2}\left(\frac{\chi_{2}^{2}}{\tilde{\omega}_{2}}-\frac{\chi_{1}^{2}}{\tilde{\omega}_{1}}\right)+\frac{\chi_{2}\delta\chi_{2}}{\Delta_{m}^{2}}-\frac{\tilde{\omega}_{2}\delta\omega_{2}}{\Delta_{m}^{2}}-\frac{\chi_{1}\delta\chi_{1}}{\Delta_{m}^{1}}$ and $\mu=(2\delta\tilde{\omega}-i\delta\Gamma_{\rm{eff}})$. In these expressions, $\delta\omega_{2}$ and $\delta\gamma$ are the two types of perturbation used in the main text. The former is induced by a mass deposition on the second mechanical resonator, while the latter comes from optical fluctuations or a mismatch damping of the two mechanical resonators. We stress that there can be also a perturbation induced by the fluctuation of the parametric driving which can be captured through $\delta\chi_{j}$ in Eq. 46. However, we have not consider such a perturbation in our investigation in the main text. We have also made some assumptions for a seek of simplicity, $\tilde{\omega}_{1}=\tilde{\omega}_{2}=\omega_{m}$ and $\chi_{1}=\chi_{2}=\chi$. With these approximations, we were able to derive the different expressions for sensitivity and enhancement factor presented in the main text, which were qualitatively in good agreement with our results. In order to give a direct comparison between the two sensing scenarios, we plotted both their sensitivity and enhancement factor in Fig. 7. The full line is from the sensing based on the shifting mechanism while the dashed line refers to splitting at the $\rm{EP}$. It can be clearly seen that the scheme based on the splitting performs better as discussed in the main text.
Figure 7: Comparison of the sensitivities and the enhancement factors for the two sensing schemes when $\chi=0$. The used parameters are those as in Fig. 7.
Appendix C Transmission and output spectra
[…]
$\mathcal{O}(\omega)=\mathcal{D}(\omega)\left[i\sqrt{\mathcal{K}}\mathcal{N}(\omega)+\sqrt{\gamma_{m}}\mathcal{O}^{in}(\omega)\right],$ (47)
where $\mathcal{D}(\omega)=(-\mathcal{M}-i\omega)^{-1}$. We are interested to obtain the spectra two optical modes at the output of the opto-mechanical cavity. According to the input-output theory, the operators for the output fields are related to the cavity and to the input noise operators by the relation $\delta\alpha^{out}_{1}(\omega)=\sqrt{\kappa}\delta\alpha_{1}(\omega)-\delta\alpha^{in}_{1}(\omega)$ and $\delta\alpha^{out}_{2}(\omega)=\sqrt{\kappa}\delta\alpha_{2}(\omega)-\delta\alpha^{in}_{2}(\omega)$. Then by using Eq.(A) and Eq.(47), the expressions for the two transmitted fields in Fourier domain are given by,
$\delta\alpha^{out}_{1}(\omega)$ $=S_{11}(\omega)\delta\alpha^{in}_{1}(\omega)+S_{12}(\omega)\delta\alpha^{in^{\dagger}}_{2}(\omega)+S_{13}(\omega)\delta\beta^{in}_{1}(\omega)$
$+S_{14}(\omega)\beta^{in}_{2}(\omega),$ (48a)
$\alpha^{out^{\dagger}}_{2}(\omega)$ $=S_{21}(\omega)\delta\alpha^{in}_{1}(\omega)+S_{22}(\omega)\delta\alpha^{in^{\dagger}}_{2}(\omega)+S_{23}(\omega)\delta\beta^{in}_{1}(\omega)$
$+S_{24}(\omega)\beta^{in}_{2}(\omega)$ (48b)
where
$S_{11}(\omega)=1-\dfrac{2\tilde{G}_{1}\xi_{2}(\omega)\sqrt{\Gamma_{1}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{12}(\omega)=-\dfrac{2i\tilde{G}_{1}\tilde{J}_{m}\sqrt{\Gamma_{2}}}{\sqrt{\kappa}\xi_{eff}(\omega)},$ (49a)
$S_{13}(\omega)=\dfrac{2i\tilde{G}_{1}\xi_{2}(\omega)\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{14}(\omega)=\dfrac{2\tilde{G}_{1}\tilde{J}_{m}\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)},$ (49b)
$S_{21}(\omega)=\dfrac{2i\tilde{G}_{2}\tilde{J}_{m}\sqrt{\Gamma_{1}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{22}(\omega)=1+\dfrac{2\tilde{G}_{2}\xi_{1}(\omega)\sqrt{\Gamma_{2}}}{\sqrt{\kappa}\xi_{eff}(\omega)},$ (49c)
$S_{23}(\omega)=\dfrac{2\tilde{G}_{2}\tilde{J}_{m}\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{24}(\omega)=-\dfrac{2i\tilde{G}_{2}\xi_{1}(\omega)\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)}$ (49d)
where $\xi_{eff}(\omega)=\xi_{1}(\omega)\xi_{2}(\omega)+\tilde{J}^{2}_{m}$ with $\xi_{1}(\omega)=\Gamma^{1}_{eff}/2+i(\Delta^{1}_{m}-\omega)$ and $\xi_{2}(\omega)=\Gamma^{2}_{eff}/2+i(\Delta^{2}_{m}-\omega)$. The input quantum noise operators cavity and mechanical modes are characterized by the following correlations
$\langle\delta\alpha^{in}_{j}(t),\delta\alpha^{in^{\dagger}}_{j}(t^{\prime})\rangle$ $=\delta(t+t^{\prime}),$ (50a)
$\langle\delta\beta^{in^{\dagger}}_{j}(t^{\prime}),\delta\beta^{in}_{j}(t)\rangle$ $=N_{s}\delta(t+t^{\prime}),$ (50b)
[…]
where $N^{j}_{s}=(n^{j}_{m}+1)\sinh^{2}(r_{j})+n^{j}_{m}\cosh^{2}(r)$ and $M^{j}_{s}=(2n^{j}_{m}+1)\sinh(r_{j})\cosh(r_{j})$ with $n^{j}_{m}=[e^{\hbar\omega_{j}/K_{B}T}-1]^{-1}$ is the mean thermal occupation number of the mechanical modes. Here we assumed that both mechanical resonators are at same temperature $T$. By using the expressions from Appendix C and Appendix C, one can derive the output spectra defined as
[…]
$=\left|S_{11}(\omega)\right|{}^{2}+\left|S_{12}(\omega)\right|{}^{2}+\left(2N_{s_{1}}+1\right)\left|S_{13}(\omega)\right|{}^{2}$
$+\left(2N_{s_{2}}+1\right)\left|S_{14}(\omega)\right|{}^{2},$ (51a)
[…]
$=\left|S_{21}(\omega)\right|{}^{2}+\left|S_{22}(\omega)\right|{}^{2}+\left(2N_{s_{1}}+1\right)\left|S_{23}(\omega)\right|{}^{2}$
$+\left(2N_{s_{2}}+1\right)\left|S_{24}(\omega)\right|{}^{2}.$ (51b)
where $N_{s_{1,2}}=(n_{m_{1,2}}+1)\sinh^{2}(r_{1,2})+n_{m_{1,2}}\cosh^{2}(r_{1,2})$.